# Aesthetics is Cheap, Show me the Text: An Empirical Evaluation of State-of-the-Art Generative Models for OCR

### v2

@@ 4.3 Handwritten Text Removal
We evaluate the handwritten text removal task, with results presented in Fig. 12. Two evaluation settings are adopted: erasing all text or erasing partial handwriting. Under the all-text removal case, given an image with simple structures (consisting of only one graphic, the text, and the background) as shown in the first row, most models can erase all handwritten text in the image. Conversely, in images with complex structures where handwritten text is embedded inside and interleaved with printed text, like the third row, they mostly fail to precisely remove the handwriting. On the other hand, as shown in the second and fourth rows, models generally fall short in removing the specific handwriting. They suffer from precisely locating the target text, mistakenly removing other text that should be preserved, and changing the structure or appearance of the background. This assessment suggests the shortage of current generative models in removing handwritten text, underscoring the necessity of future dedicated optimization.
5 Result of Scene Text Image
@@
Figure 13: T2I generation results for scene text images. The red glowing indicates the relatively best output corresponding to each input. We test three language cases: English, Chinese, and mixed-language. Most models can handle English scene text generation. Yet, only GPT-4o and Qwen-VLo-preview handle Chinese text generation, Flux.1-Kontext-dev even delivers a totally unrelated result. For mixed-language, all models fail to handle this case, demonstrating chaotic, blurred, and nonsensical generated text. Even the relatively optimal-performing GPT-4o fails to generate all texts accurately, particularly small ones. For complete prompts, please refer to our GitHub repository.
@@
Figure 14: Text editing and text removal results for scene text images. In scene text editing, while GPT-4o successfully modifies the target text, it adds unexpected text (the first row) and compromises the original aspect ratio. Other models exhibit inadequate text modification precision and substantial loss of textual and background texture fidelity. In scene text removal, models only fulfill the requirements in the cases of the first row, while accidentally eliminating other details like the animal and auxiliary text, or failing to erase smaller texts. This means scene text removal remains a challenge for current generative models.
@@
Scene text, or text in the wild, refers to textual information that appears naturally in real-world environments. It can be found on product packaging, vehicle license plates, street signs, and numerous other objects in our daily surroundings. Unlike documents or handwriting, scene text is often characterized by its unconstrained font appearances and diverse backgrounds, such as curved text [105], multiple text orientations [106], perspective distortion [107], and low contrast and cluttered background [108]. Recent research has expanded beyond detection and recognition [96, 109, 110, 111] to focus on generative tasks [69, 62, 61, 112] involving scene text. These include T2I generation for synthesizing images with specific text content, scene text editing for modifying existing text while maintaining visual coherence, and scene text removal for erasing text while preserving background integrity. These generative applications support content creation, visual design, and privacy protection, though they remain challenging due to the complex integration of text within visual contexts. Therefore, we evaluate existing models on various scene text generation tasks to benchmark their capabilities in handling the complexity and diversity of text in real-world scenarios.
@@
Figure 15: T2I generation results for images with naturally embedded text. As observed in the first and second rows, GPT-4o is pretty good at generating images with a small amount of text, perfectly following the instructions, while other models either fail to generate specified text or synthesize blurred, unclear text. However, in the third and fourth rows, all models can not accurately generate ordered text like numbers on the ruler or letters on the keyboard. This may be attributed to their lack of world knowledge or the insufficiency in synergizing their innate knowledge and generation performance. For complete prompts, please refer to our GitHub repository.
5.1 T2I Generation
@@
The T2I generation result for scene text images is shown in Fig. 13. We present three cases featuring English, Chinese, and mixed-language text to comprehensively demonstrate the models’ scene text generation capabilities across different languages. Most models can accurately generate English scene text, with only BAGEL and Janus-4o failing in this case. However, only GPT-4o and Qwen-VLo-Preview can handle the Chinese scene text generation, while Flux.1-Kontext-dev produces a completely irrelevant image without any text, suggesting it may lack Chinese prompt understanding capability. For mixed languages, none of the models can accurately generate all the text, with most producing chaotic, blurred, or non-semantic content. We observe that although Qwen-VLo-Preview performs better in generating Chinese and English scene text, it struggles with other languages in the multilingual case, indicating potential limitations in this area, while GPT-4o demonstrates more balanced capabilities.
5.2 Scene Text Editing and Scene Text Removal
@@
We also evaluate these models’ text editing and text removal capabilities on scene text images, as presented in Fig. 14. For text editing, only GPT-4o can accurately modify text according to instructions, but it also adds unexpected text (the first row) and fails to maintain the original aspect ratio of the input image. Other models fail to accurately modify the text and lose other textual or background texture details. For text removal, although some models can erase target text, many details remain problematic, such as changing background textures, failing to erase smaller text, or removing text and background that should be preserved. Moreover, we observed two main issues with these generative models when performing these two tasks: (1) Instruction understanding problems, where Flux.1-Kontext-dev cannot understand Chinese instructions and preserves the original image. (2) Text generation problems, where Janus-4o performs poorly in both modifying text and preserving the original text. In conclusion, scene text editing and scene text removal remain challenging for current models.
5.3 Naturally Embedded Text
Naturally embedded text refers to text that appears as an integral part of objects or scenes in images, where the text is seamlessly incorporated into the physical elements (e.g., smartwatch, keyboard) rather than being artificially overlaid. It is also a type of scene text, but differs from conventional scene text such as street signs and billboards, typically exhibiting greater diversity and complexity in terms of font, size, orientation, material, and integration methods. These texts often need to adapt to the functional requirements and physical constraints of the devices they appear on, such as display resolution, ergonomic keyboard design, or precise scale requirements of measuring tools, resulting in visual characteristics and recognition challenges distinctly different from ordinary scene text. Therefore, we specifically evaluate existing models’ T2I generation and text editing capabilities on naturally embedded text.
5.3.1 T2I Generation
@@
We demonstrate the models’ T2I generation capabilities for naturally embedded text in Fig. 15. As observed in the first and second rows, GPT-4o performs excellently when generating electronic display fonts, and other models either fail to generate specified text or synthesize blurred, unclear text. However, when asked to generate items with ordered text, such as rulers or keyboards, all models struggle significantly; despite most being able to generate the correct items, they all exhibit problems with inaccurate ruler measurements and unreasonable keyboard characters. For instance, the rulers often show inconsistent or illogical measurement intervals and values, while the keyboards frequently display scrambled or repetitive characters instead of the standard QWERTY layout. This may be attributed to their lack of world knowledge or the insufficiency in synergizing their innate knowledge and generation performance, particularly in cases requiring precise spatial arrangement and sequential ordering of text elements.
@@
Figure 16: Text editing results for images with naturally embedded text. Text editing in these cases involves fine-grained and small character modifications. From the results, GPT-4o performs best, successfully editing most texts to their target contents. However, it sometimes fails to preserve original text or pixel texture details (the second and third rows). Other models primarily struggle to modify small text (such as changing “5 km” to “63 km”) and sometimes output the vanilla input unchanged (OmniGen2 and BAGEL). Therefore, small text editing remains a significant challenge for existing generative models.
5.3.2 Text Editing
@@
Finally, we evaluate the text editing task on naturally embedded text, with results shown in Fig. 16. We provide instructions for models to attempt to modify fine-grained and small text in images. Among all tested models, GPT-4o is the best-performing one, only failing when modifying Chinese text (the third row), though it lacks the ability to maintain the original text’s texture details. Other models primarily struggle to modify much smaller text (like changing "5 km" to "63 km") and sometimes directly output the vanilla input without changes (OmniGen2 and BAGEL). These results indicate that small text editing remains a significant challenge for existing generative models, particularly in cross-language application scenarios and maintaining visual consistency, providing clear directions for future optimization of multimodal models.
6 Result of Artistic Text Image
Artistic text is stylized typography that incorporates creative graphical components or decorative fonts to achieve aesthetic appeal beyond standard readable formatting. It has found widespread applications from creating memorable brand logos and advertisements to designing impactful movie titles, game interfaces, and personalized merchandise [113]. In the recent decade, numerous synthesis methods [114, 115, 116, 117] have been proposed to assist automatic glyph design, demonstrating remarkable font fidelity and diversity. Therefore, we incorporate this task into our evaluation, testing whether current generative models can fulfill this interesting, creative, and practical task.
6.1 T2I Generation
We first assess the T2I generation task, with results presented in Fig. 17. For line-level artistic text generation, the results are indeed inferior. Although existing models produce creative, visually pleasing artistic fonts, they suffer from content missing in both English and Chinese cases, which is not feasible for real-world applications. In addition, we impose on them to generate complex, rare Chinese characters, and they consistently fail to fulfill the requirements. The Flux.1-Kontext-dev stands as the worst performer, which generates human and photo frame images instead of text images. For character-level generation, all models can generate English single-character. However, for the rare Chinese character, they have failed again.
@@ 6.1 T2I Generation
Figure 17: T2I generation results for artistic images. The outcomes of line-level text generation are inferior. All models, despite the relatively better GPT-4o, can not completely and accurately render long text in the artistic fonts. Also, they fail to generate images of rare Chinese characters. For complete prompts, please refer to our GitHub repository.
@@
Figure 18: Text editing and style transfer results for artistic images. Overall, the editing and style transfer of the English script are well done by most models. Conversely, for the Chinese language, GPT-4o is the only model that can handle Chinese text editing and style transfer, which, however, still accidentally squares the image. This indicates that apart from the English language, multilingual text modification abilities need further improvement for current models.
6.2 Text Editing and Style Transfer
@@
Subsequently, we assess the text editing and style transfer abilities of current generative models. Results are shown in Fig. 18. Here, style transfer looks similar to text editing due to their shared objectives of modifying the source text to the target. Hence, it is necessary to clarify the definitions of these two tasks: text editing requires the models to modify the text and keep other elements of the image unchanged, while style transfer aims at generating a new image with specific text content according to the source image’s text style, regardless of other elements in the vanilla images.
@@
We evaluate text editing under two scenarios: including editing the text w/o and w/ style modifications. In the first and third rows, we instruct the model to modify text content while preserving the original style. All models demonstrate strong performance on English text, although GPT-4o, Qwen-VLo-Preview, and Janus-4o produce square outputs instead of maintaining the original rectangular aspect ratio. For Chinese text editing, only GPT-4o and Qwen-VLo-Preview show capability, but with notable artifacts: GPT-4o alters the background to black, while Qwen-VLo-Preview introduces unnecessary decorative elements. In the second row, most models successfully modifies both text content and writing style, while preserving background elements like fireworks and cityscape views. A minor issue is that some models turn the original rectangular images into squares.
@@
As for style transfer, the source style are well transferred to the target content in the English cases. However, similar to text editing, most models can not handle style transfer of Chinese text. Collectively, these findings indicate that current generative models can handle English artistic text editing and style transfer effectively, but substantial improvements are needed for Chinese script processing.
7 Result of Complex and Layout-Rich Text Image
@@
Complex and layout-rich text images, such as slides and posters, blend text, graphics, and intricate layouts to convey information effectively and aesthetically, which are widely used in marketing, education, and entertainment. The dynamic and context-sensitive nature of these layouts requires models to comprehend spatial arrangements, hierarchy, and visual balance. For instance, in a slide, text should align seamlessly with graphics to emphasize key points without causing clutter [118, 119, 120]. In addition, posters demand bold typography and striking visuals to capture attention [121, 122, 123, 124, 125]. Successfully generating and editing such images demands not only an understanding of textual semantics but also proficiency in design principles.
@@
Figure 19: T2I generation and text editing results for slide images. GPT-4o demonstrates significantly better instruction-following abilities and generation quality in T2I generation than other models, which fulfills most requirements. However, other models fail to generate exact English/Chinese text but in unknown languages, also generating square images even given the instruction of generating a retangular one. Regarding text editing, actually all models fall short in this case, given rich and dense text like slide images. Locating fine-grained and a small portion of text to perform modification proves a great challenge for current generative models.
@@
Figure 20: T2I generation and text editing results for poster images. Most models can handle the generation of English posters, whereas they fall short in generating Chinese posters, primarily attributed to their lack of Chinese training data. Similar to Fig. 19, current models mostly cannot accurately modify text, and even when they can modify text, they cannot guarantee that other original details remain unchanged.
@@
Figure 21: Results of the layout-aware text generation task. For English cases, OmniGen2 is the only one that adds correct text in an appropriate position without obstructing the original main component and preserving other pixel details. For Chinese cases, GPT-4o stands as the only model that adds correct text in appropriate positions. Other models seem to lack the ability to comprehend and generate Chinese text.
7.1 Slide Image
7.1.1 T2I Generation and Text Editing
@@
Typically, slide images consist of rich and structured elements such as headings, text blocks, and graphics, ordered in a hierarchical layout. This presents great challenges to generation models for comprehending both textual semantics and design aesthetics in slide creation. The generated results are presented in the top frame of Fig. 19. As observed, while GPT-4o suffers from text errors, it demonstrates significantly better instruction-following abilities than other models, producing readable text and more appealing layouts. In contrast, other models generate illegible text and sometimes output totally unrelated images. For example, Qwen-VLo-Preview produces an image of handwritten text but a slide. Similarly, Flux.1-Kontext-dev fails to interpret Chinese instructions, instead generating an unrelated image of a young girl. These results highlight the challenges generative models face in producing complex slide content.

### v3

@@ 4.3. Handwritten Text Removal
In addition, we evaluate the handwritten text removal task, with results presented in Fig. 12. Two evaluation settings are adopted: erasing all text or erasing partial handwriting. Under the all-text removal scenario, given an image with simple structures, (i.e., comprising a single graphic, text, and background as shown in the first row), most models successfully erase all handwritten text within, except Qwen-Image, which produces two puzzling words. Conversely, when presented images with complex structures where handwritten text is embedded inside and interleaved with printed text, like the third row, most of them fail to precisely remove the handwriting. For partial removal, as shown in the second and fourth rows, models generally perform poorly. They often fail to precisely locate the target text, mistakenly remove other text that should be preserved, and alter the structure or appearance of the background. These results highlight the shortage of current generative models in removing handwritten text, underscoring the necessity of dedicated optimization.
5. Result of Scene Text Image
@@
Figure 13. T2I generation results for scene text images. We test three language cases: English, Chinese, and mixed-language. Most models can handle English scene text generation. Yet, only GPT-4o and Qwen-VLo-preview handle Chinese text generation; Flux.1-Kontext-dev even delivers a totally unrelated result. For mixed-language, only Qwen-Image delivers a nearly perfect result, followed by GPT-4o and Qwen-VLo-Preview. Other models fail to handle this case, demonstrating chaotic, blurred, and nonsensical generated text. For complete prompts, please refer to our GitHub repository.
@@
Figure 14. Text editing and text removal results for scene text images. In scene text editing, while GPT-4o successfully modifies the target text, it adds unexpected text (the first row) and alters the original aspect ratio. Other models exhibit inadequate text modification precision and substantial loss of textual and background texture fidelity. In scene text removal, models only fulfill the requirements in the first row case, while accidentally eliminating other details like the animal and auxiliary text, or failing to erase smaller texts.
@@
Scene text, or text in the wild, refers to textual information that appears naturally in real-world environments, including vehicle license plates, street signs, and numerous other objects in daily surroundings. Unlike documents or handwriting, scene text features unconstrained font appearances and diverse backgrounds, such as curved text (Liu et al., 2020a), multiple text orientations (Zhou et al., 2017), perspective distortion (Shi et al., 2019), and low contrast and cluttered background (Lee and Osindero, 2016). Recent research has expanded beyond detection and recognition (Luo et al., 2019; Huang et al., 2022; Liu et al., 2023; Huang et al., 2024) to encompass generative tasks involving scene text (Liu et al., 2020b; Tuo et al., 2024; Chen et al., 2023; Peng et al., 2024), including scene text synthesis (T2I generation), scene text editing, and scene text removal. Therefore, we evaluate existing models across these three tasks to benchmark their capabilities in handling the complexity and diversity of text in real-world scenarios.
@@
Figure 15. T2I generation results for images with naturally embedded text. In the first and second rows, the first three models perfectly follow the instructions and output applicable results, while other models either fail to generate specified text or synthesize blurred, unclear text. However, in the third and fourth rows, all models can not accurately generate ordered text like numbers on the ruler or letters on the keyboard. For complete prompts, please refer to our GitHub repository.
5.1. T2I Generation
@@
Fig. 13 presents T2I generation result for scene text images. We evaluate three cases that feature English, Chinese, and mixed-language text to comprehensively demonstrate the models’ scene text generation capabilities across different languages. Most models can accurately generate English scene text, with only BAGEL and Janus-4o failing in this case. However, only GPT-4o, Qwen-VLo-Preview, and Qwen-Image can handle the Chinese scene text generation, while Flux.1-Kontext-dev produces a completely irrelevant image without any text, suggesting it may lack Chinese prompt understanding capability. For mixed languages, no model can accurately generate the text, though Qwen-Image produces a nearly perfect outcome (only “Bazaar” is wrongly rendered as “Baaar”). Other models mainly produce chaotic, blurred, or non-semantic content. In addition, we observe that Qwen-VLo-Preview struggles with other languages in the multilingual case, while GPT-4o and Qwen-Image deliver more balanced performance across diverse languages.
5.2. Scene Text Editing and Scene Text Removal
@@
We also evaluate these models’ text editing and text removal capabilities on scene text images, as presented in Fig. 14. For text editing, only GPT-4o accurately modifies text according to instructions, but it still adds unexpected text (the first row) and does not preserve the original aspect ratio of the input image. The other models fail to modify the text and lose other textual or background texture details. For text removal, although some models can erase target text, many details remain problematic, such as changing background textures, failing to erase smaller text, or removing text and background that should be preserved. On both instruction-based editing tasks, Qwen-Image and Janus-4o demonstrate the worst performance: the former misunderstands instructions and edits the wrong text and the latter produces ruined images. These results indicate that scene text editing and removal remain unaddressed challenges for current generative models.
5.3. Naturally Embedded Text
Naturally embedded text refers to text seamlessly incorporated into physical elements (e.g., smartwatch, keyboard) rather than artificially overlaid. Due to its natural integrity, we define them as a type of scene text with different materials and integration methods. Hence, we specifically evaluate existing models’ T2I generation and text editing capabilities on naturally embedded text.
5.3.1. T2I Generation
@@
We demonstrate the models’ T2I generation capabilities for naturally embedded text in Fig. 15. As observed in the first and second rows, the first three models perform excellently when generating electronic display fonts, and other models either fail to generate specified text or synthesize blurred, unclear text. However, when asked to generate items with ordered text, such as rulers or keyboards, all models struggle significantly. For instance, the rulers often show inconsistent or illogical measurement intervals and values, while the keyboards frequently display scrambled or repetitive characters instead of the standard QWERTY layout. This may be attributed to their lack of world knowledge or the insufficiency in synergizing their innate knowledge and generation performance, particularly in cases requiring precise spatial arrangement and sequential ordering of text elements.
@@
Figure 16. Text editing results for images with naturally embedded text. GPT-4o successfully edits most texts but occasionally fails to preserve some original text or pixel texture details (the second and third rows). Other models primarily struggle to modify small text and directly output vanilla images or irrelevant results.
5.3.2. Text Editing
@@
Furthermore, we evaluate text editing on naturally embedded text, with results shown in Fig. 16. We mainly instruct models to modify small text in images. Among all tested models, GPT-4o performs best, though it fails when modifying Chinese text (third row) and falls short in maintaining pixel texture details. Other models primarily struggle to modify much smaller text (like changing “5 km” to “63 km”) and sometimes directly output the vanilla input unchanged (OmniGen2 and BAGEL). Qwen-Image even generates a human picture given a fridge panel input. The results indicate that small text editing remains a significant challenge for existing generative models, particularly in cross-language applications and visual consistency maintenance. This provides clear directions for future optimization of multimodal models.
6. Result of Artistic Text Image
@@
Figure 17. T2I generation results for artistic images. Only Qwen-Image generates complete content, while other models fail to completely and accurately render text lines with artistic fonts. Also, all models fail to generate images of rare Chinese characters. For complete prompts, please refer to our GitHub repository.
Artistic text is stylized typography that incorporates creative graphical components or decorative fonts to achieve aesthetic appeal beyond standard readable formatting. It has found widespread applications from creating memorable brand logos and advertisements to designing eye-catching movie titles, game interfaces, and personalized merchandise (Bai et al., 2024). In the recent decade, numerous synthesis methods (Gao et al., 2019; Wang et al., 2023; Li et al., 2023b; Tanveer et al., 2023) have been proposed to assist automatic glyph design, demonstrating remarkable font fidelity and diversity. Therefore, we incorporate this task into our evaluation, testing whether current generative models can fulfill this interesting, creative, and practical task.
@@
Figure 18. Text editing and style transfer results for artistic images. Most models perform well on English script editing and style transfer. However, for Chinese text, only GPT-4o can handle editing and style transfer, though it still inadvertently crops the image. This indicates that current models’ multilingual text modification abilities require significant improvement beyond English.
6.1. T2I Generation
We first assess the T2I generation task, with results presented in Fig. 17. For line-level text generation, Qwen-Image is the only model that generates correct content in simplified Chinese and English scenarios. While other models produce visually pleasing artistic fonts, they suffer from content missing. In addition, Additionally, when tasked with generating complex, rare Chinese characters, they consistently fail to meet requirements. Flux.1-Kontext-dev performs worst, which generates human and photo frame images instead of text images. For character-level generation, all models can generate English single-character. However, they again fail to generate rare Chinese characters.
6.2. Text Editing and Style Transfer
@@
Subsequently, we evaluate the text editing and style transfer abilities of current generative models. Results are shown in Fig. 18. Here, style transfer looks similar to text editing due to their shared objectives of modifying the source text to the target. Hence, it is necessary to clarify the definitions of these two tasks: text editing requires the models to modify the text and keep other elements of the image unchanged, while style transfer aims at generating a new image with specific text content according to the source image’s text style, regardless of other elements in the vanilla images.
@@
We evaluate text editing under two scenarios: editing the text without and with style modifications. In the first and second rows, we instruct models to modify text while preserving the original style. Most models demonstrate strong performance on English text. For Chinese text, only GPT-4o and Qwen-VLo-Preview can perform the task, although both introduce artifacts: GPT-4o alters the background to black, while Qwen-VLo-Preview introduces unnecessary decorative elements. In the third row, most models successfully modify both text content and writing style, while preserving background elements like fireworks and cityscapes. Notably, Qwen-Image adds the new text without replacing the original, and models like GPT-4o and Janus-4o can turn the original rectangular images into squares, which is an unexpected minor issue.
@@
For style transfer, the source style is well transferred to the target content in English cases. However, similar to text editing, most models can not handle style transfer of Chinese text. Collectively, these findings indicate that current models can handle English artistic text editing and style transfer effectively but require further improvements for Chinese script processing.
7. Result of Complex & Layout-Rich Text Image
@@
Complex and layout-rich text images, such as slides and posters, blend text, graphics, and intricate layouts for both visual aesthetics and effective communication, which are widely used in marketing, education, and entertainment. While automated generation of these images grows increasingly popular, it requires understanding both textual semantics and design principles. For instance, slides require text seamlessly aligned with graphics to emphasize key points without clutter (Fu et al., 2022; Bandyopadhyay et al., 2024; Zheng et al., 2025), while posters demand bold typography and striking visuals to capture attention (Pang et al., 2025; Chen et al., 2025c; Peng et al., 2025; Gao et al., 2025). This drives us to include this type of text image in the evaluation.
@@
Figure 19. T2I generation and text editing results for slide images. GPT-4o and Qwen-Image outperform other models in instruction-following and generation quality, meeting most requirements. However, other models generate unknown languages when English/Chinese is requested, also yielding square images despite instructions for rectangular ones. Regarding text editing, all models struggle with dense slide text. This proves that locating and modifying a small portion of text remains a great challenge for current generative models.
@@
Figure 20. T2I generation and text editing results for poster images. Most models can handle the generation of English posters but fall short in generating Chinese posters, primarily attributed to their lack of Chinese training data. Qwen-Image delivers the only successful Chinese poster generation case. Similar to Fig. 19, models largely fail at accurate text editing, and even when edits succeed, original details are often altered.
@@
Figure 21. Layout-aware text generation results for CLR text. For English cases, only OmniGen2 and GPT-4o add correct text in the proper position without occluding the main component, also preserving other details. For Chinese cases, only GPT-4o succeeds, while other models struggle to comprehend or generate Chinese text.
7.1. Slide Image
7.1.1. T2I Generation
@@
We first investigate the slide generation abilities of existing models, with results presented in the top frame of Fig. 19. As observed, GPT-4o and Qwen-Image are the top two performers, which demonstrate stronger instruction-following abilities and text rendering abilities than other models, despite some content errors. In contrast, other models generate illegible text or even totally unrelated images. For example, Qwen-VLo-Preview produces an image of handwritten text but a slide. Similarly, Flux.1-Kontext-dev fails to interpret Chinese instructions, instead generating an unrelated image of a young girl. These results highlight the challenges generative models face in producing complex slide content.
7.1.2. Text Editing